# Canonical formalism in the Schwarzschild spacetime

### v2

@@ IV.3.1 Derivation
The spin actions $\mathcal{J}_{i}$ in Eqs. (137) are a generalization of geodesic actions defined in Refs. [46, 53] to the spinning setting, through linear order in the body’s spin; see also Refs. [65, 108] for other attempts to define action variables for a spinning body. These spin actions are the functions of constants of motion $P_{\alpha}$ in Eq. (134). Because the two sets $P_{\alpha}$ and $\mathcal{J}_{i}$ are bijective by definition, we immediately find that
@@ IV.4.2 Spin dynamics
In particular, applying the relationship between the spin angle ${\varpi}$ and the Marck (precession) angle $\psi_{p}$ established in Eq. (132) to the circular orbit demonstrates that the solution in Eq. (153) is identical to the Marck precession angles $\psi_{p}$ (modulo the offset phase $\phi_{s}$) derived in Refs. [42, 43, 36]; see, e.g., Eqs. (4.25) and (4.26) in Ref. [42] (in Schwarzschild limit) 242424 The minus sign on the right-hand side of Eq. (153) is a consequence of the right-handed tetrad in our formalism, in contrast to the left-handed Marck tetrad that is typically employed in the literature (cf. Refs. [35, 42, 43]). A more detailed discussion of this distinction can be found in Appendix C.1.. This result is a direct consequence of the fact that the angle $\Theta$ in Eq. (132) can be set to zero for circular orbits, given that we have $\cos\Theta=0$ (and $\sin\Theta=1$) when $\tilde{\pi}_{r}=0$.
@@ IV.4.3 ISCO parameters
All the circular and ISCO values of the parameters $\rho,\varepsilon,l$ derived above were first analytically computed for equatorial orbits and spin-aligned configurations in a Schwarzschild spacetime in Appendix B.3 of [109] (see also Refs. [110, 111, 112, 113] for earlier examinations of the ISCO of a spinning body in Kerr spacetime.), with all subsequent literature [cf. Refs. [114, 42, 43, 40] agreeing with them. We have verified that our formulae are in perfect agreement with their results presented in Eqs. (B17) through (B20b) in Ref. [109] [^25] There is, however, a typo in Eq. (B.18) of Ref. [109], where the denominator of the second term in the right-hand side should read $r-3m_{2}$ instead of $r-2m_{2}$..
@@ Acknowledgements.
Appendix A Conventions and notations
[…]
The symbol $O(\epsilon^{N})$ indicates an expression of order $N$ in the small parameter $\epsilon=S_{\circ}/\mu M$ [cf. Eq. (14)]. Unless otherwise specified, we work consistently at order $O(\epsilon)$ and neglect any $O(\epsilon^{n})$ with $n\geq 2$. Different sets of coordinate systems used in the body of the paper are summarized in Table. 2.
Table 2: List of different coordinate system used in this work:
Manifold Coordinate set     Canonical Definition
14D-Poisson $(x^{\alpha},p_{\beta},S^{\gamma\delta})$ No Paper I
manifold $\mathcal{M}$ $y_{\mathcal{M}}:=(x^{\alpha},\pi_{\beta},S^{I},D^{I})$ No Paper I
12D-Symplectic $y_{\mathcal{N}}:=(t,r,\theta,\phi,\sigma,\zeta,\pi_{t},\pi_{r},\pi_{\theta},\pi_{\phi},\pi_{\sigma},\pi_{\zeta})$ Yes Paper I
manifold $\mathcal{N}$ $y_{\mathcal{N}}^{\textrm{A}}:=(t,r,\nu,\ell,s,\zeta,\pi_{t},\pi_{r},\pi_{\nu},\pi_{\ell},\pi_{s},\pi_{\zeta})$ Yes Andoyer
10D-Physical $y_{\mathcal{P}}^{\textrm{C}}:=(t,r,\nu,\ell,\Gamma,\pi_{t},\pi_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi})$ No Complete
phase space $\mathcal{P}$ $y_{\mathcal{P}}^{\textrm{F}}:=(\tilde{t},r,\nu,\tilde{\ell},{\varpi},\pi_{t},\tilde{\pi}_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi})$ Yes Final
Appendix B The non-spinning (geodesic) Hamiltonian
@@
B.1 Schwarzschild spherical symmetry
@@
B.2 The orbital plane as a $1:1$ resonance
@@
Appendix C TD spin vectors and parallel transport
@@
The methods summarized above originate from Marck [34], are widely implemented in literature. Details in the case of Kerr spacetime are provided in Refs. [34, 115, 35, 31], and we review the main points of Marck’s method here; see also Sec. III.D of Ref. [42]. In this appendix, we work only with the geodesic orbit $\mathscr{L}_{0}$ in Schwarzschild spacetime, where spin effects to the orbital dynamics are negligible in Eq. (175), and omit the “g” label to unclutter the notation.
C.1 The Marck tetrad
A way to obtain the orthonormal tetrad is to make use of the solution to the geodesic orbit $\mathscr{L}_{0}$ (though different from the tetrad in Eq. (33) adopted in our Hamiltonian formulation). Their explicit construction and resulting expressions in Kerr spacetime are displayed in, e.g., Sec. 3 of Ref. [34] and Eqs.(47) - (51) in Ref. [35]. We specialize these results to Schwarzschild spacetime with the Schwarzschild-Droste coordinate $x^{\alpha}=(t,\,r,\,\theta,\,\phi)$ and four constants of motion $(\mu,E,L_{z},L)$ of $\mathscr{L}_{0}$ [cf. Appendix. B]. Following the tradition in the literature, we use the assignment [it is understood that $(r,\,\theta)$ = $(r({\bar{\tau}}),\,\theta({\bar{\tau}}))$ below]
@@ C.1 The Marck tetrad
C.2 Parallel and perpendicular components
@@
Appendix D Details relative to the reduction from $\mathcal{N}$ to $\mathcal{P}$
@@
D.1 Computing Poisson brackets involving the TD SSCs
@@
D.2 Fixing the homogeneous freedom in the transformation in Eq. (106)

### v3

@@ IV.3.1 Derivation
The spin actions $\mathcal{J}_{i}$ in Eqs. (134) are a generalization of geodesic actions defined in Refs. [68, 75] to the spinning setting, through linear order in the body’s spin; see also Refs. [66, 57, 140, 94, 95] for other attempts to define action variables for a spinning body. These spin actions are the functions of constants of motion $P_{\alpha}$ in Eq. (131). Because the two sets $P_{\alpha}$ and $\mathcal{J}_{i}$ are bijective by definition, we immediately find that
@@ IV.4.1 Orbital parameters
All the circular values of the parameters $\varepsilon,l$ derived above were first analytically computed for equatorial orbits and spin-aligned configurations in a Schwarzschild spacetime;see, e.g., in Appendix B.3 of [141]. These findings have since been corroborated by other authors [cf. Refs. [142, 63, 64, 61].]
@@ IV.4.2 Spin dynamics
In particular, applying the relationship between the spin angle ${\varpi}$ and the Marck (precession) angle $\psi_{p}$ established in Eq. (186) to the circular orbit demonstrates that the solution in Eq. (150) is identical to the Marck precession angles $\psi_{p}$ (modulo the offset phase $\phi_{s}$) derived in Refs. [63, 64, 57]; see, e.g., Eqs. (4.25) and (4.26) in Ref. [63] (in Schwarzschild limit) 232323 The minus sign on the right-hand side of Eq. (150) is a consequence of the right-handed tetrad in our formalism, in contrast to the left-handed Marck tetrad that is typically employed in the literature [cf. Refs. [56, 63, 64, 61]]. A more detailed discussion of this distinction can be found in Appendix B.1.. This result is a direct consequence of the fact that the angle $\Theta$ in Eq. (186) can be set to zero for circular orbits, given that we have $\cos\Theta=0$ (and $\sin\Theta=1$) when $\tilde{\pi}_{r}=0$.
@@ IV.4.3 ISCO parameters
We have verified that our ISCO formulae obtained in this section are all in perfect agreement with the results presented in Eqs. (B17) through (B20b) in Ref. [141]  [^24] There is, however, a typo in Eq. (B.18) of Ref. [141], where the denominator of the second term in the right-hand side should read $r-3m_{2}$ instead of $r-2m_{2}$.; given that their small parameter (denoted $s$ there) is our $-K/\pi_{\phi}$. See also Refs. [143, 144, 145, 146] for earlier examinations of the ISCO of a spinning body in Kerr spacetime.
@@ Acknowledgements.
Appendix A The non-spinning (geodesic) Hamiltonian
@@
A.1 Schwarzschild spherical symmetry
@@
A.2 The orbital plane as a $1:1$ resonance
@@
Appendix B TD spin vectors and parallel transport
@@
The methods summarized above originate from Marck [55], are widely implemented in literature. Details in the case of Kerr spacetime are provided in Refs. [55, 147, 56, 51], and we review the main points of Marck’s method here; see also Sec. III.D of Ref. [63]. In this appendix, we work only with the geodesic orbit $\mathscr{L}_{0}$ in Schwarzschild spacetime, where spin effects to the orbital dynamics are negligible in Eq. (172), and omit the “g” label to unclutter the notation  [^26] The difference between the true orbit $\mathscr{L}$ and $\mathscr{L}_{0}$ is negligible here because $S_{a}^{\text{\tiny{TD}}}$ is already a quantity of $O(\epsilon)$..
B.1 The Marck tetrad
A way to obtain the orthonormal tetrad is to make use of the solution to the geodesic orbit $\mathscr{L}_{0}$ Their explicit construction and resulting expressions in Kerr spacetime are displayed in, e.g., Sec. 3 of Ref. [55] and Eqs.(47) - (51) in Ref. [56]. We specialize these results to Schwarzschild spacetime with the Schwarzschild-Droste coordinate $x^{\alpha}=(t,\,r,\,\theta,\,\phi)$ and four constants of motion $(\mu,E,L_{z},L)$ of $\mathscr{L}_{0}$ [cf. Appendix A]. It is helpful to note that the Marck tetrad is distinct from the natural-basis tetrad introduced in Eq. (34) and utilized in our Hamiltonian formulation.
Following the tradition in the literature, we use the assignment [it is understood that $(r,\,\theta)$ = $(r({\bar{\tau}}),\,\theta({\bar{\tau}}))$ below]
@@ B.1 The Marck tetrad
B.2 Parallel and perpendicular components
@@
B.3 Comparison with the canonical representation
The TD spins $S_{a}^{\text{\tiny{TD}}}$, as expressed in Eq. (180) in terms of the Marck tetrad, can be directly compared with an alternative expression for $S_{a}^{\text{\tiny{TD}}}(\tilde{x}^{i},\tilde{\pi}_{i})$ in terms of the canonical coordinates $(\tilde{x}^{i},\tilde{\pi}_{i})$ in Eq. (115) in the phase space $\mathcal{P}$. This allows us to gain more insight into the physical interpretation of the spin canonical variables $({\varpi},\,\pi_{\varpi})$ [cf. Eqs. (115)] in the context of Schwarzschild spacetime.
The relevant Marck’s expressions to the comparison with Eq. (129) and (130) are  [^28] Notice that we here adopt the right-handed representation of the Marck tetrad, instead of the left-handed one widely used in the literature [cf. Refs. [66, 56, 51, 63]]; the details are explained in Appendix. B.1.
and
$p_{\phi}S_{\theta}^{\text{\tiny{TD}}}-p_{\theta}S_{\phi}^{\text{\tiny{TD}}}=-\frac{rL}{\csc\theta}{\mathcal{S}}_{\parallel}\,.$ (182)
where we used $\pi_{\alpha}=p_{\alpha}+O(\epsilon)$. Here, the constant-angle $\phi_{s}$ defines the orientation of the TD spin (in the Marck frame), and ${\psi_{p}}={\psi_{p}}({\bar{\tau}})$ describes the Marck’s precession phase that grows with the proper time $\bar{\tau}$ according to the differential equation in Eq. (177).
Equation (182) immediately allows us to explain why the constant of motion $\pi_{\varpi}$ in Eq. (119) is interpreted as the “parallel component” of the spin. A direct comparison of these equations demonstrates that
$\pi_{\varpi}={\mathcal{S}}_{\parallel}\,,\quad\text{with}\quad S_{\circ}^{2}-\pi_{\varpi}^{2}={\mathcal{S}}_{\perp}\,,$ (183)
and we find that $\pi_{\varpi}$ is equivalent to the amplitude of the third leg of Marck tetrad ${\mathcal{S}}_{\parallel}$ aligned with the direction of the Killing-Yano angular momentum vector ${\cal L}^{a}$. Similarly, the overall amplitude $\sqrt{S_{\circ}^{2}-\pi_{\varpi}^{2}}$ in Eq. (129) can be interpreted as the “perpendicular” component of the TD spin ${\mathcal{S}}_{\perp}=\sqrt{S_{\circ}^{2}-{\mathcal{S}}_{\parallel}^{2}}$, in the sense that it is perpendicular to ${\cal L}^{a}$.
Upon inserting Eq. (183) within Eq. (129), $S_{t}^{\text{\tiny{TD}}}(\tilde{x}^{i},\tilde{\pi}_{i})$ and $S_{r}^{\text{\tiny{TD}}}(\tilde{x}^{i},\tilde{\pi}_{i})$ can be directly compared with each corresponding component in Eq. (181) through the covariance of the TD spin $S^{a}_{\text{\tiny{TD}}}$ (in contrast to the non-covariant $3$-spin vectors $\vec{S}$). This enables a comparison of the phases $\varpi$ and $\phi_{s}+\psi_{p}$ observed in the TD spin. Using $p_{a}=\pi_{a}+O(\epsilon)$ and $L=\pi_{\ell}$, this comparison gives rise to two linear relations between the phases of the TD spin, which are expressed in the matrix form
[…]
It is in this sense that the variable ${\varpi}$ is related to the “spin precession”. When the canonical variable ${\varpi}$ is rotated by the angle $\Theta$ (in the phase space $\mathcal{P}$), it coincides with the Marck precession phase $\psi_{p}$ with the phase offset $\phi_{s}$ (in the Schwarzschild spacetime).
As a conclusion of this subsection, we need to comment on the angle $\Theta$ that appears in Eq. (186). The physical meaning of this angle is currently unknown. We have attempted to derive Eqs. (185) and  (187) within our Hamiltonian formalism (i.e. without relying on the comparison with Marck’s representation of $S_{a}^{\text{\tiny{TD}}}$), but, for reasons that remain unclear to us, this approach does not seem to not seem to be effective in this case.292929 Technically, it was challenging to infer the specific combination $r^{2}\mu^{2}+\pi_{\ell}$ that enters the denominator of Eq. (185) (as well as Eq. (177)) from the canonical Hamiltonian system. The question of why the Marck’s precession phase $\psi_{p}$ must be a superposition of the canonical variable ${\varpi}$ and ”the” angle $\Theta$ (when viewed from the phase space $\mathcal{P}$) is a profound one; we leave it open for the time being.
Appendix C Details relative to the reduction from $\mathcal{N}$ to $\mathcal{P}$
@@
C.1 Computing Poisson brackets involving the TD SSCs
@@
C.2 Fixing the homogeneous freedom in the transformation in Eq. (109)